# thanks: Brinson Prize Fellow

### v2

@@ VI.3 Measurement of the subhalo mass function in CDM
Restricting to regions of parameter space with $m_{\rm{hm}}<10^{5}\mathrm{M}_{\odot}$ removes most of the effects of WDM free streaming in the model-predicted datasets and flux ratio statistics. This prior therefore takes the CDM limit of the joint posterior distribution, which we can use to infer the amplitude of the differential subhalo mass function in CDM (Equation 15). In Figure 20, we show the posterior distribution $p\left(\Sigma_{\rm{sub}}|\bf{D}\right)$. At $95\%$ confidence we infer $\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}=-0.6_{-0.5}^{+0.4}$, a more precise measurement of subhalo abundance relative to omitting imaging data $\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}=-0.8_{-0.9}^{+0.5}$.
@@
Here, we assume a typical host halo mass $10^{13}\mathrm{M}_{\odot}$ for a deflector at $z_{\rm{d}}=0.5$, such that $\mathcal{F}\left(m_{\rm{host}},z_{\rm{d}}\right)=1$ in Equation 16. Given the width of the prior on $\alpha$, and the fact that the posterior on $\alpha$ is not constrained within the range of this prior, we set $\alpha=-1.9$. Our inference used $m_{H}=10^{10.7}\mathrm{M}_{\odot}$ and $m_{L}=10^{6}\mathrm{M}_{\odot}$, respectively, with a pivot scale $m_{0}=10^{8}\mathrm{M}_{\odot}$, and our tidal stripping model predicts a mean bound mass fraction $\bar{f}_{\rm{bound}}=0.05$ in CDM [Du++25]. With these numbers we find $\Sigma=6\times 10^{7}\left(\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\right)$. Our inference on $\Sigma_{\rm{sub}}$ therefore corresponds to $\Sigma=1.6_{-1.1}^{+2.4}\times 10^{7}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$ at $95\%$ confidence. This is statistically consistent with the predictions from galacticus, $\Sigma=0.6\times 10^{7}\mathrm{M}_{\odot}/\rm{kpc^{-2}}$ [Gannon++25, Du++25], but in mild tension with the predictions from recent $N$-body simulations presented by Nadler++23, $\Sigma=0.3\times 10^{7}\mathrm{M}_{\odot}/\rm{kpc^{-2}}$.
@@
Figure 20: The posterior distribution of $\Sigma_{\rm{sub}}$ obtained from taking the CDM limit through a prior on the half-mode mass $\pi\left(\log_{10}m_{\rm{hm}}/\mathrm{M}_{\odot}\right)=\mathcal{N}\left(4.0,0.5\right)$. The color scheme is the same as Figure 19. Vertical dashed lines represent $95\%$ confidence intervals. Predictions from the semi-analytic model $\tt{galacticus}$ and $N$-body simulations are marked with vertical dotted lines.
For $f_{\rm{DM}}=0.5$ [Auger++10], and $\Sigma_{\rm{crit}}=2\times 10^{9}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$, the critical density for lensing assuming typical lens and source redshifts $z_{\rm{d}}=0.5$ and $z_{\rm{s}}=2.0$, Equation 27 gives $f_{\rm{sub}}=0.012\left(\Sigma_{\rm{sub}}/0.1\ \rm{kpc^{-2}}\right)$. Our inference on $\Sigma_{\rm{sub}}$ using imaging data therefore corresponds to projected mass fraction bound to dark matter subhalos $f_{\rm{sub}}=3.2_{-2.2}^{+4.9}\%$ at $95\%$ confidence.
@@ VII Discussion and conclusions
  • Assuming a log-uniform prior on the normalization of the subhalo mass function $\Sigma_{\rm{sub}}$ that spans a factor of 100 in subhalo abundance, incorporating constraints from lensed arcs improves bounds on dark matter warmth by 0.8 dex, leading to $m_{\rm{hm}}<10^{8.5}\mathrm{M}_{\odot}$ (Bayes factor 10:1). These weaker constraints relative to some previous work are the result of extending the $\Sigma_{\rm{sub}}$ prior to higher values, which reveals a significant covariance between $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$. Breaking this covariance with a prior on $\Sigma_{\rm{sub}}$ based on the semi-analytic model galacticus and $N$-body simulations results in $m_{\rm{hm}}<10^{7.4}\mathrm{M}_{\odot}$ and $m_{\rm{hm}}<10^{7.2}\mathrm{M}_{\odot}$, respectively. These bounds are stronger than those obtained without imaging data by 0.5 and 0.3 dex, respectively. The half-mode mass upper limits of $10^{7.4}M_{\odot}$ and $10^{7.2}M_{\odot}$ correspond to lower bounds on the mass of a thermal relic dark matter particle $7.4$ keV and $8.4$ keV, respectively.
  • Assuming CDM, we infer a projected surface mass density in subhalos $1.6_{-1.1}^{+2.4}\times 10^{7}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$, and a projected mass fraction near the typical Einstein radius of a strong lens system $f_{\rm{sub}}=3.2_{-2.2}^{+4.9}\%$. These measurements of subhalo abundance are statistically consistent with the predictions of the semi-analytic model galacticus, but in mild $\sim 2\sigma$ tension (higher) than the predictions from recent $N$-body simulations.
@@ VII.2 Interpretation
Our constraints on the free-streaming length are stronger than those obtained from other individual small-scale probes, including previous lensing studies [Birrer++17, Gilman++20, Hsueh++20, Keeley++24], the Lyman-$\alpha$ forest [Villasenor++23, Irsic++24], dwarf galaxies [Newton++21, Nadler++21, Dekker++22, Tan++25] and stellar streams [Banik++21, Nibauer++25], although probe combination between lensing and dwarf galaxies has resulted in stronger constraints [Nadler++21]. Our limits are to be considered robust, as the numerous improvements to our modeling framework overall weaken the bounds on WDM for a given dataset with respect to previous analyses. Our measurement of the projected mass density in CDM subhalos is consistent with the predictions from the semi-analytic model galacticus to within $2\sigma$ [Gannon++25, Du++25]. It is in mild tension (inferring a higher value) with the predictions from the recent Symphony suite of $N$-body simulations [Nadler++23, Gannon++25].
We note that galacticus was calibrated using the caterpillar suite of $N$-body simulations [Griffen++16], which predict more subhalos than the Symphony $N$-body simulations [Nadler++23]. The origin of this difference is not yet clearly understood. The higher subhalo abundance predicted by semi-analytic models could suggest that recent $N$-body simulations still suffer from artificial subhalo disruption, or that halo finders do not locate all subhalos in the inner regions of the host, despite recent improvements on this front [e.g. Mansfield++24] that reveal more substructure than previously reported.
@@ VII.3 Discussion and outlook
Extending our analysis to a larger sample of lenses would reduce the statistical uncertainties of the measurement, and push constraints on substructure properties to even lower halo mass scales. This will soon be possible, as cosmological surveys, such as Euclid, the Rubin Observatory, and the Roman Space Telescope discover thousands of new strong lens systems [OguriMarshall10, Shajib++25]. Higher quality imaging data, particularly for systems that lack HST imaging, or where the lensed arcs in HST imaging have low signal to noise, would enable stronger constraints on the mass profile of the main deflector, leading to more robust dark matter inferences. As discussed in Section V.5, in several systems we discovered satellite galaxies of the main deflector by visual inspection of HST data. In other cases, we use high-resolution imaging to identify stellar disks in the main deflector, and omit these systems from the sample. These considerations underscore the importance of having high angular resolution space-based imaging of the main deflector.

### v3

@@ VI.3 Measurement of the subhalo mass function in CDM
Restricting to regions of parameter space with $m_{\rm{hm}}<10^{5}\mathrm{M}_{\odot}$ removes most of the effects of WDM free streaming in the model-predicted datasets and flux ratio statistics. This prior therefore takes the CDM limit of the joint posterior distribution, which we can use to infer the amplitude of the differential subhalo mass function in CDM (Equation 15). In Figure 20, we show the posterior distribution $p\left(\Sigma_{\rm{sub}}|\bf{D}\right)$. At $95\%$ confidence we infer $\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}=-0.57_{-0.55}^{+0.42}$, a more precise measurement of subhalo abundance relative to omitting imaging data $\log_{10}\Sigma_{\rm{sub}}/\rm{kpc^{-2}}=-0.76_{-0.95}^{+0.46}$. As discussed in Section IV.3, to convert these values to the amplitude of the bound mass function in CDM, for which tidal stripping is approximately independent of subhalo infall mass, one can use Equation 15 with a normalization at $10^{8}M_{\odot}$ given by $\bar{f}_{\rm{bound}}\times\Sigma_{\rm{sub}}$, with $\bar{f}_{\rm{bound}}=0.05$ [Du++25].
@@
Here, we assume a typical host halo mass $10^{13}\mathrm{M}_{\odot}$ for a deflector at $z_{\rm{d}}=0.5$, such that $\mathcal{F}\left(m_{\rm{host}},z_{\rm{d}}\right)=1$ in Equation 16. Given the width of the prior on $\alpha$, and the fact that the posterior on $\alpha$ is not constrained within the range of this prior, we set $\alpha=-1.9$. Our inference used $m_{H}=10^{10.7}\mathrm{M}_{\odot}$ and $m_{L}=10^{6}\mathrm{M}_{\odot}$, respectively, with a pivot scale $m_{0}=10^{8}\mathrm{M}_{\odot}$. With these numbers we find $\Sigma=6.2\times 10^{7}\left(\Sigma_{\rm{sub}}/\rm{kpc^{-2}}\right)$. Our inference on $\Sigma_{\rm{sub}}$ therefore corresponds to $\Sigma=1.7_{-1.2}^{+2.6}\times 10^{7}\ \mathrm{M}_{\odot}\rm{kpc^{-2}}$ at $95\%$ confidence. This is statistically consistent with the predictions from galacticus, $\Sigma=0.9\times 10^{7}\ \mathrm{M}_{\odot}\rm{kpc^{-2}}$. It is in mild $\sim 1.8\sigma$ tension with the predictions from the recent Symphony $N$-body simulations presented by Nadler++23, $\Sigma=0.6\times 10^{7}\mathrm{M}_{\odot}\rm{kpc^{-2}}$. These predictions from $N$-body simulations and galacticus follow from the theoretical priors on $\Sigma_{\rm{sub}}$ discussed in Section IV.3.
@@
Figure 20: The posterior distribution of $\Sigma_{\rm{sub}}$ obtained from taking the CDM limit through a prior on the half-mode mass $\pi\left(\log_{10}m_{\rm{hm}}/\mathrm{M}_{\odot}\right)=\mathcal{N}\left(4.0,0.5\right)$. The color scheme is the same as the left panel of Figure 19. Vertical dashed lines represent $95\%$ confidence intervals. Predictions from the semi-analytic model $\tt{galacticus}$ and $N$-body simulations (see discussion in Section IV.3) are marked with vertical lines.
For $f_{\rm{DM}}=0.5$ [Auger++10], and $\Sigma_{\rm{crit}}=2.0\times 10^{9}\mathrm{M}_{\odot}\ \rm{kpc^{-2}}$, the critical density for lensing assuming typical lens and source redshifts $z_{\rm{d}}=0.5$ and $z_{\rm{s}}=2.0$, Equation 27 gives $f_{\rm{sub}}=0.012\left(\Sigma_{\rm{sub}}/0.1\ \rm{kpc^{-2}}\right)$. Our inference on $\Sigma_{\rm{sub}}$ using imaging data therefore corresponds to a projected mass fraction bound to dark matter subhalos $f_{\rm{sub}}=3.2_{-2.3}^{+5.1}\%$ at $95\%$ confidence.
@@ VII Discussion and conclusions
  • Assuming a log-uniform prior on the normalization of the subhalo mass function $\Sigma_{\rm{sub}}$ that spans a factor of 100 in subhalo abundance, incorporating constraints from lensed arcs improves bounds on dark matter warmth by 0.6 dex, leading to $m_{\rm{hm}}<10^{8.6}\mathrm{M}_{\odot}$ (Bayes factor 10:1). These weaker constraints relative to some previous work are the result of extending the $\Sigma_{\rm{sub}}$ prior to higher values, which reveals a significant covariance between $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$. Breaking this covariance with a prior on $\Sigma_{\rm{sub}}$ based on the semi-analytic model galacticus and $N$-body simulations results in $m_{\rm{hm}}<10^{7.4}\mathrm{M}_{\odot}$ and $m_{\rm{hm}}<10^{7.2}\mathrm{M}_{\odot}$, respectively. These bounds are stronger than those obtained without imaging data by 0.4 dex. The half-mode mass upper limits of $10^{7.4}M_{\odot}$ and $10^{7.2}M_{\odot}$ correspond to lower bounds on the mass of a thermal relic dark matter particle $7.4$ keV and $8.4$ keV, respectively, using Equation 18.
  • Assuming CDM, we infer a projected surface mass density in subhalos $1.7_{-1.2}^{+2.6}\times 10^{7}\ \mathrm{M}_{\odot}\rm{kpc^{-2}}$, and a projected mass fraction near the typical Einstein radius of a strong lens system $f_{\rm{sub}}=3.2_{-2.3}^{+5.1}\%$. These measurements of subhalo abundance are statistically consistent with the predictions of the semi-analytic model galacticus, but in mild tension (higher) than the predictions from recent $N$-body simulations.
@@ VII.2 Interpretation
Our constraints on the free-streaming length are stronger than those obtained from other individual small-scale probes, including previous lensing studies [Birrer++17, Gilman++20, Hsueh++20, Keeley++24], the Lyman-$\alpha$ forest [Villasenor++23, Irsic++24], dwarf galaxies [Newton++21, Nadler++21, Dekker++22, Tan++25] and stellar streams [Banik++21, Nibauer++25], although probe combination between lensing and dwarf galaxies has resulted in stronger constraints [Nadler++21] by breaking covariance between $\Sigma_{\rm{sub}}$ and $m_{\rm{hm}}$. Our limits are to be considered robust, as the numerous improvements to our modeling framework overall weaken the bounds on WDM for a given dataset with respect to previous analyses.
Our measurement of the projected mass density in CDM subhalos is consistent with the predictions from the semi-analytic model galacticus, but in mild tension (inferring a higher value) with the predictions from the Symphony suite of $N$-body simulations [Nadler++23]. The differences between subhalo abundance predicted by semi-analytic models and $N$-body simulations could suggest that recent $N$-body simulations still suffer from artificial subhalo disruption, or that halo finders do not locate all subhalos in the inner regions of the host, despite recent improvements on this front [e.g. Mansfield++24] that reveal more substructure than previously reported.
@@ VII.3 Discussion and outlook
Extending our analysis to a larger sample of lenses would reduce the statistical uncertainties of the measurement, and push constraints on substructure properties to even lower halo mass scales. This will soon be possible, as cosmological surveys, such as Euclid, the Rubin Observatory, and the Roman Space Telescope discover thousands of new strong lens systems [OguriMarshall10, Shajib++25]. Higher quality imaging data, particularly for systems that lack HST imaging, or where the lensed arcs in HST imaging have low signal to noise, would enable stronger constraints on the mass profile of the main deflector, leading to more robust dark matter inferences. As discussed in Section V.5, in several systems we discovered satellite galaxies of the main deflector by visual inspection of HST data. In other cases, we use high-resolution imaging to identify stellar disks in the main deflector, and omit these systems from the sample. These considerations underscore the importance of having high angular resolution space-based or adaptive optics imaging of the main deflector.